### v1

# Searching for internal symbols underlying deep learning

Abstract
Deep learning (DL) enables deep neural networks (DNNs) to automatically learn complex tasks or rules from given examples without instructions or guiding principles. As we do not engineer DNNs’ functions, it is extremely difficult to diagnose their decisions, and multiple lines of studies proposed to explain principles of DNNs/DL operations. Notably, one line of studies suggests that DNNs may learn concepts, the high level features recognizable to humans. Thus, we hypothesized that DNNs develop abstract codes, not necessarily recognizable to humans, which can be used to augment DNNs’ decision-making. To address this hypothesis, we combined foundation segmentation models and unsupervised learning to extract internal codes and identify potential use of abstract codes to make DL’s decision-making more reliable and safer.
1 Introduction
Deep learning (DL) can automatically wire deep neural networks (DNNs) to perform complex tasks, and DNNs’ learning capability allows them to outperform humans in some tasks [1, 2, 3]. Notably, instead of expert instructions or feature engineering, DL requires only training examples to solve highly complex problems. However, as DNNs’ structures are automatically determined without guiding principles, we do not fully understand their decision-making process. That is, DNNs may be useful black box decision-makers, but without safety measures, deploying them in high-stakes domains (e.g., autonomous cars) comes with unpredictable risks [4, 5]. For the moment, the current generation of DL does not warrant essential safety to be deployed in safety-critical or high-stakes domains. To address this shortcoming, earlier studies sought potential methods that could decode and explain DNNs’ decisions. There are multiple lines of research in explainable AI (e.g., see Refs. [6, 7, 8, 9, 10]), but the most relevant to this study is the one probing correlations between hidden layer representations and inputs or between hidden layer representations and outputs (i.e., decisions). Feature visualization [11] and network dissection [12] sought direct links between inputs and hidden layers, while linear probes investigated correlations between outputs and hidden layers [13, 14].
It is notable that the mapping between hidden layers and visual features in input layers, in principle, establishes links between high level patterns recognizable to humans (often referred to as ‘concepts’) and hidden neurons. The Test with Concept Activation Vectors [14] explicitly extended this idea to detect concepts encoded in neural networks. The concepts (i.e., perceptible patterns) coded in neural networks could deepen our understanding of DNNs’ operating principles and eventually provide explanations (readable to humans) of DNNs’ decisions. More importantly, it raises the possibility that DNNs could build some abstract codes in hidden layers, some of which could directly be linked to concepts. In this study, we extended this line of studies by seeking abstract codes, which are not necessarily recognizable to humans, in hidden layers functionally linked to DNNs’ decisions. We also explored potential methods of utilizing hidden layers’ abstract codes to make DNNs’ operations/decision-making more reliable and safer.
To make a clear distinction between human recognizable concepts and DNNs’ own abstract codes, we will refer to DNNs’ abstract codes as ‘symbols’ hereafter. Our analysis of hidden layer representations is based on four assumptions. First, symbols ($S_{i}$, abstract codes of DNNs) are encoded across feature maps in each layer. Second, responses of individual hidden layers $H_{i}$ can be described by the sum of symbols $S_{i}$ and residuals $\epsilon$; $H_{i}=S_{i}+\epsilon$. Conversely, we can map individual representations $H_{i}$ into nearest $S_{i}$ by minimizing $\epsilon$. Third, there is a finite number of symbols in DNNs, which is far less than the possible number of hidden layer representations. Fourth, symbols are spatially local and thus need to be extracted from proper receptive fields or regions of interest (ROIs). With these assumptions in mind, we performed unsupervised clustering analysis with hidden layer responses ($H_{i}$) to determine the cluster centers and identify ‘symbols’. Our analysis suggests that these symbols can mediate semantic meanings and allow us to 1) monitor the models’ decision-making process, 2) detect abnormal operations related to adversarial perturbations and out-of-distribution (OOD) examples and 3) temporarily learn about OOD examples. Based on the results, we argue that hidden layers’ responses can be converted to discrete symbols that can help us build more reliable and safer DNNs.
2 Extracting symbols underlying DNNs’ operations
To test if DNNs learn to use symbols to make decisions, we analyzed ImageNet models’ hidden layer responses to a subset (Mixed_13) of ImageNet [15], previously curated by a python machine learning library [16]. Mixed_13 contains 78 classes out of 1,000 classes of ImageNet, which belong to one of 13 super-classes (Supplementary Table1). Mixed_13 consists of two disjointed ‘training’ and ‘test’ sets. To extract the symbols, we assumed that the symbols associated with DNN’s decision-making would be observed frequently. For instance, if a symbol is associated with a tiger, a picture of a tiger would elicit it frequently. With this assumption in mind, we presented the inputs from Mixed_13 and searched for repeatedly occurring hidden layer responses by using unsupervised clustering analysis.
We note that analyzing hidden layer responses can be challenging for two reasons. First, hidden layer responses are extremely noisy high-dimensional data. A massive number of hidden neurons exist in modern DNNs, and their responses are determined with diverse inputs, making dimension reduction essential (see [17, 18, 13] , for instance). Second, visual scenes often contain multiple objects rather than a single one. For instance, a visual scene may include a tiger or tigers walking along the river, playing with other tigers and/or climbing a tree. Thus, if we want to extract codes related to tigers, it is necessary to isolate regions of interest (ROIs) containing tigers and exclude confounding objects (e.g., tree or river). However, ImageNet training examples are labeled with single class names and do not provide sufficient information to identify ROIs.
To address these challenges, we first used segmentation models to extract ROIs and recorded hidden layer responses spatially corresponding to ROIs. Then, we used ROI-pooling operation, proposed for object detectors [19], to extract smooth 3-by-3 activation vectors per single feature map. Once the activation vectors were extracted, we conducted clustering algorithms to determine the codes repeatedly occurring with Mixed_13 (a subset of ImageNets). In section 2.1, we discuss how ROIs are extracted from Mixed_13. In section 2.2, we present the details of extracting hidden layer responses. In section 2.3, we explain our clustering analysis in detail.
2.1 ROIs Identification
We note that ‘Second Thought Certification’ (STCert) proposed in an earlier study [20] can be used to identify ROIs related to visual objects with high accuracy. STCert is a two-step process (Fig. 1). First, it uses foundational segmentation models [21, 22] to detect ROIs related to DNNs’ predictions. For instance, if a DNN predicts a tiger, STCert asks the segmentation model to return the bounding box (i.e., ROI) of the tiger in the picture. Second, it crops ROIs and uses them to confirm or reject DNNs’ original predictions. STCert provides a more reliable ROI estimates than a single step segmentation, since its segmentation is crosschecked with a classifier. As such, we used STCert to identify ROIs in this study to extract ROIs from randomly chosen 200 training examples for each class in Mixed_13; see Supplementary text A.1 for more detail. Since Mixed_13 contains 78 classes (Supplementary Table1), the total number of training examples is 15,600. Among them, we first selected examples, on which DNNs’ predictions were correct to minimize the bias induced by the incorrect predictions/decisions of DNNs (i.e., incorrect mapping between ROI and the inputs’ labels). For some of the images, STCert did not find ROIs or found multiple ROIs, which means the actual number of ROIs are determined by a few factors, not just the number of training examples.
Figure 1: Schematics of STCert [20].
2.2 Generating hidden layer activity vectors
Once ROIs were identified, we sampled activations ($H_{i}$) of hidden neurons, whose receptive fields correspond to ROIs. To reduce noise and filter out non-essential information, we used ROI-pooling [19] to convert hidden layer activations into 3-by-3 activation maps. That is, the size of hidden layer activation maps can be arbitrary depending on the size of objects, but ROI-pooling always turns them into 3-by-3 activation maps. It should be noted that individual activations in 3-by-3 activation maps reflect pixels at different spatial locations. As we assumed that symbols are collectively encoded neurons that share the same receptive field, we built vector codes (referred to as activity vectors below) by collecting activations at the same location in the maps across channels. For instance, there are 512 channels in ResNet18’s layer 4, and we have 9 location-specific activation maps in each channel. In this case, we aggregated activations from the same location in the activation map from all 512 channels and then created 9 activity vectors, each of which contains 512 components, from the layer per image.
In this study, we analyzed 4 hidden layers of popular DNNs, ResNet18, ResNet50, VGG19 and DenseNet121 [23, 24, 25]. ResNet and DenseNet consist of 4 composite blocks, and their outputs were analyzed. VGG19 contains 5 max pooling layers, whose outputs roughly correspond to ResNet’s composite block’s outputs. We selected the last 4 max-pooling layers and analyzed them, as they match the layers of ResNet and DenseNet. We used Pytorch [26] and Timm [27] python deep learning libraries to implement these models. A workstation with Intel Core9 CPU and RTX 4090 was used for all our experiments and analyses.
2.3 Clustering analysis of hidden layer activity vectors
We turned to unsupervised clustering analysis to identify representative activity vectors equivalent to symbols in this study. We also note that many activity vectors contain low values. With these low-value activity vectors, clustering algorithms consider the zero-valued vector as the most prominent cluster center, which is not informative. Thus, we removed the vectors consisting of values lower than the mean activity value estimated on each layer. Next, we used UMAP analysis [28, 29] to further reduce them to 3-dimensional vectors; see Table 1 for the parameters used in the analysis. These three dimensional vectors were analyzed to detect symbols via clustering. It should be noted that the exact number of clusters existing in the activity vector space are unknown, and the result of clustering analysis strongly depends on the number of predefined clusters. To address this issue, we used X-means clustering [30] to automatically discover an optimal number of clusters (i.e., the number of symbols). We note that X-means is an extension of Kmeans clustering, which uses Bayesian Information Criterion (BIC) to determine the optimal number of cluster, and a python package ‘pyclustering’ [31] was used to implement X-means in this study. In our analysis, we restrict the maximum number of classes to 1,000. Table 2 shows the identified number of clusters in 4 layers of DNNs tested. As shown in the table, the optimal numbers of clusters in early layers are smaller than 1,000, but those in the late layers hit the maximum value, suggesting that homogeneity of activity vectors increased, as layers go deeper.
Table 1: List of Parameters used in UMAP analysis.
Parameter Value
n_neighbors 50
min_dist 0.1
n_component 3
metric Euclidean
3 Symbols Analysis
3.1 Links between symbols and semantic meanings of inputs
Table 2: Number of symbols identified from 4 models. For VGG19, we list 4 last max-pooling layers, and for all other layers, we list 4 outputs of 4 functional blocks.
Layer ResNet18 ResNet50 VGG19 DenseNet121
1 960 967 837 895
2 940 992 1000 1000
3 1000 1000 1000 1000
4 1000 1000 1000 1000
Then, we asked if the identified symbols (i.e., cluster centers) can be linked to semantic meanings of inputs (i.e., labels) by correlating them with inputs’ labels. Specifically, for each image ($img_{k}$), we obtained ROIs ($ROI_{m}$) and 9 symbols ($S_{mn}$, where $n=1,...,9$). These 9 symbols $S_{mn}$ were correlated with the labels ($l_{m}$) of $ROI_{m}$, which were determined by the label of $img_{k}$ containing $ROI_{m}$. That is, Correlation Map ($CM(i,j)$) evolves over all identified ROIs obtained from the training examples of Mixed_13 according to Eq. 1.
$CM(i,j)=\begin{cases}CM(i,j)+1&\text{if $i=S_{mn}$, $j=l_{m}$},\text{where }n=1,...,9\\ CM(i,j)&\text{otherwise}\par\end{cases}$ (1)
, where $CM(i,j)$ is the component of Correlation map at $i^{th}$ row and $j^{th}$ column; $S_{mn}$ denotes the $n^{th}$ (out of 9) symbol related to $m^{th}$ ROI; where $l_{m}$ is the label of $m^{th}$ ROI (i.e., $img_{k}$).
If a symbol is associated with a specific class, it would appear frequently when the class object is presented, but it would not appear when other class objects are presented. Fig. 2 shows the correlation $CM(i,j)$ between 100 randomly chosen symbols from 4 layers of ResNet18 and 78 classes. $Y$-axis denotes 100 randomly chosen symbols, and $x$-axis denotes the label (i.e., class). We found two types of symbols. First, some symbols were associated with a broad range of examples. Second, some symbols were selectively linked to specific classes. Symbols specific to classes were rare in early layers but became more common in late layers. We found equivalent results in ResNet50, VGG19 and DenseNet121 (Supplementary Fig. 1). In addition, we note that many symbols can be merged together (Supplementary Fig. 2) and thus, did not increase the maximum number of clusters above 1000. These results suggest that some symbols could be directly linked to inputs’ labels, and thus, one could infer the labels from their associated symbols.
Figure 2: The correlations between the symbols of ResNet18 and classes. We show the correlation maps $CM(i,j)$ between 100 symbols randomly chosen and all 78 classes of Mixed_13. $y$-axis denotes the indices of symbols, $x$-axis denotes the class. (A)-(D), the correlations observed in layers 1-4, respectively.
To test this possibility, we obtained the symbols associated with test examples of Mixed_13 and asked if they can be used to predict the labels of inputs (i.e., ROIs). In this experiment, we collected ROIs from all test images whether or not DNNs make correct predictions. From these ROIs, we collected activity vectors and obtained the test symbols (i.e., the nearest cluster centers). These ‘test symbols’ were used to predict the likelihood of each class by using $CM(i,j)$ (Eq. 1). As we aimed to compute the probability of the class, the correlation map $CM(i,j)$ was normalized along with a row (i.e., over classes) using SoftMax (2). By using the normalized Correlation Map $P(i,j)$, we calculated the probability $P(i,j)^{m}$ that $m^{th}$ ROI belonged to the class $j$ when the symbol $S_{i}$ was observed (Eq. 2). That is, symbol indices were used as hash keys for a look-up table.
$P(i=S_{i},j)^{m}=\frac{e^{C(i=S_{i},j)}}{\sum_{n=1}^{78}e^{C(i=S_{i},n)}}$ (2)
As we obtained 9 symbols from each ROI, we took average probability over 9 (Eq. 3) and then determine the most probable class as the prediction $Pr_{m}$ (Eq. 4).
$P(j)^{m}=\frac{1}{9}\Sigma_{i=1}^{9}P(i,j)^{m}$ (3)
$Pr_{m}=\operatorname*{argmax}_{j}P(j)^{m}$ (4)
Fig. 3 shows the accuracy of the symbol-based predictions depending on the layer. We note that the prediction accuracy is around 80% in layer 4 and much lower in earlier layers. These accuracies cannot be directly compared to the models’ accuracy because the choice is limited to one out of 78 classes, not 1000 classes. However, the accuracy in layer 4 and earlier layers were significantly higher than chance (1/78), supporting that the identified symbols were associated with semantic meanings.
If symbols are associated with semantic meanings, they could also be associated with DNNs’ predictions. We tested this possibility by correlating the symbols and DNNs’ predictions on Mixed_13 training examples (once again, 200 examples per class) instead of inputs’ labels. As stated above, we used the symbols to infer their predictions on test examples and found that the symbols are predictive of DNNs’ answers on test examples (Fig. 3B), raising the possibility that DNNs could utilize symbols to make decisions. The observed link between symbols and DNNs’ decisions leads us to hypothesize that symbols can be used to augment DNNs’ decision-making process. Below we propose potential ways to make DNNs’ operations more reliable using symbols.
Figure 3: Symbol-based predictions. (A), Symbol-based prediction on the labels of test examples of Mixed_13. (B), Symbol-based prediction on the DNNs’ answers.
3.2 How confident are you, DNNs?
In the analysis above, we used $CM(i,j)$ to infer the most probable class by using symbol as a hash key and $CM(i,j)$ as a look-up table. However, this look-up table can be read in opposite direction. If we know the label $j$ of inputs, we can infer how likely individual symbols $S_{i}$ appear in DNNs. With this idea in mind, we defined ‘Expected Symbol Score (ESS)’ per layer with respective label/class $j$ (Eq. 5).
$ESS_{j}=\frac{1}{9}\Sigma_{S_{1},...,S_{9}}P(S_{i},j)^{m}$ (5)
, where $S_{i}$ denotes the indices of 9 symbols obtained from each ROI. That is, $ESS_{j}$ with respective class $j$ is the mean value of 9 probability from $P(i,j)^{m}$; see Eq. (5). We estimated $ESS_{j}$ from all 4 layers. First, we asked if ESSs obtained from layer 4 can be correlated with the accuracy of DNNs’ answers. Specifically, we split ESSs into two distributions depending on whether or not DNNs made the correct decisions.
We found that ESSs were significantly higher when predictions were correct (see data marked by * in Fig. 4A) than when they were incorrect (see data without * in Fig. 4A). To quantify the separation between correct and incorrect predictions, we estimated AUROC (the orange line in Fig. 4B), suggesting that ESSs can indicate DNNs’ likelihood of making correct predictions. We also examined ESSs from layers 1, 2 and 3 and AUROC between correct and incorrect answers (Fig. 4B). Further, we examined the norm of ESSs $||(ESS_{1},ESS_{2},ESS_{3},ESS_{4})||$ from all layers, and AUROC estimated from this norm were similar to those obtained from layer 4 (see blue line Fig. 4B).
Figure 4: Symbol-based confidence level. (A), Expected symbol score (ESS) in ResNet18(Res18), ResNet50(Res50), VGG19 and DenseNet121(DENSE). The asterisks denote ESSs when DNNs make correct predictions, which are higher than those when DNNs make incorrect predictions. (B), AUROC between ESS when DNNs are correct and ESS when DNNs are wrong. We tested ESS from layer 4(orange), layer 3(green), layer2(red) and layer1(violet). Also we tested the norm from all layers, and AUROC between the norms is shown in blue.
3.3 Detecting out-of-distribution (OOD) examples
Detecting OOD examples has been a central challenge in computer vision models. During training, models are forced to choose one of the predefined choices. After training, they must choose the most probable answer on any given input, even when the inputs are not related to its training. By nature, predictions on OOD examples are random and should be ignored. To address this shortcoming, OOD detection algorithms have been proposed to inform DL models (and its users) when to ignore inputs; see [32] for a review.
We assumed that hidden layer responses (i.e., symbols), highly optimized for in-distribution examples during training, are in synchronization across layers during inference, when inputs are drawn from in-distribution, but they can be out of synchronization with each other, when inputs are drawn from out-of-distribution. Further, we speculated that ESSs can be used to measure the synchrony between layers and thus detect OOD examples. To test this line of hypotheses, we extracted the symbols from NINCO dataset [33], which contain OOD examples carefully curated against ImageNet.
In our analysis, we first used $CM(i,j)$ of layer 4 to predict the most probable class, which were incorrect, as the inputs were drawn from OOD, and tested ESSs (once again, expected symbol scores) in layers 1,2 and 3. If the symbols are truly out of synchronization due to OOD examples, we anticipated low expected symbol scores in all three layers. We chose 18 classes from NINCO dataset, and ROIs of NINCO images were determined by GroundingDINO (an open-set segmentation model[21]) using their real class labels. We evaluated the expected symbol scores and compared them with the ESSs of normal images from the test set of Mixed_13.
Fig. 5A visualizes ESSs ($ESS_{1},ESS_{2},ESS_{3}$) from layers 1, 2, 3 in a 3D space. Blue circles and orange triangles denote in-distribution and OOD examples respectively. As shown in the figure, ESSs were significantly different between the in-distribution and OOD examples, and we quantified this observation by estimating AUROC. The estimated AUROC from ESSs of all 3 layers and their norms $||(ESS_{1},ESS_{2},ESS_{3})||$ suggest that ESSs are well separated between normal and OOD examples and can be used to detect OOD examples.
Figure 5: (A), Comparison of ESSs between in-distribution and OOD examples. Orange triangles denote ESSs from layers 1,2 and 3 when inputs are drawn from OOD. Blue circles denote ESSs when inputs are drawn from in-distribution examples. (B), AUROC between ESSs for all 4 DNNs. AUROC were calculated using ESSs from layers 3, 2 and 1. Additionally, ESSs were also calculated using the norm of ESS from all 3 layers.
3.4 Detecting adversarial examples
Since the discovery of DL models’ vulnerability to adversarial examples [34, 35], adversarial perturbation has emerged as one of the greatest threats to DL’s safety. Given that all inputs are progressively processed in modern DNNs, it seems natural to assume that adversarial perturbation may change internal representation on layer-by-layer basis. An earlier study [36] found that the difference in representations in early layers was not noticeable, but in late layers it grew significantly. We asked if ESSs can be used to detect the influence on adversarial perturbation. Specifically, we crafted adversarial examples from validation images of Mixed_13 using AutoAttack [37], which is known to be one of the most effective attacks. In the experiment, we used a ‘standard’ mode with $eps=0.03$ and $norm=Linf$. For individual inputs, we used labels of clean images to detect ROIs and estimated the corresponding symbols; that is, ROIs are identified with original labels, not manipulated predictions.
For ROIs, we estimated the norm of ESSs from all 4 layers $||(ESS_{1},ESS_{2},ESS_{3},ESS_{4})||$ and calculated AUROC by comparing the norms of ESSs between normal and adversarial inputs. We note that the AUROC estimated from the norms were close to 0.5 (see the blue line in Fig. 6), suggesting that the symbols in DNNs were not significantly different between normal and adversarial examples. Consistent with this finding, the symbols of adversarial inputs in layer 4 predict the original classes as accurately as the symbols of the normal inputs do (the thick green line in Fig. 6). These results suggest that the symbols are much more resilient than the models’ final predictions. That is, the detection of adversarial inputs may require methods other than comparing internal symbols.
So far, ESSs have been computed using the most probable class predicted by the layer 4 correlation map $CM(i,j)$, but they can also be computed using DNNs’ predictions. If ESSs are computed with DNNs’ predictions, we can effectively evaluate the consistency between models’ decisions and internal symbols, which we assume could be useful in detecting adversarial perturbation. However, we analyzed unmodified ImageNet models, whose output nodes correspond to 1,000 classes, not 78 of Mixed_13, Consequently, the predictions were not necessarily confined to 78 classes of Mixed_13. Thus, to build ESSs using their predictions, we selected adversarial and normal symbols when predictions are one of 78 classes (see Table 3 for the number of ROIs included in this analysis). This Decision-based ESSs are compared in Fig. 6 (see the orange line). As shown in the figure, Decision-based ESSs are significantly different for all 4 models, suggesting that ESSs with respect to DNNs’ (manipulated) predictions can be used to detect adversarial perturbation.
Figure 6: Symbol-based detection of adversarial perturbations. The orange line denotes AUROC calculated from decision-based ESS, whereas the blue dashed line denotes AUROC calculated ESS. The green line denotes the accuracy of symbol-based predictions on the adversarial inputs.
Table 3: Number of normal and adversarial ROIs. For both normal and adversarial inputs, we first selected inputs (i.e., images), on which DNNs’ predictions were one of 78 classes of Mixed_13. Then, we selected all ROIs from these examples (both normal and adversarial). We list the number of ROIs selected below. It should be noted that images are not chosen using the correctness of DNNs’ predictions.
Model Normal Adversarial
ResNet18 4002 1161
ResNet50 4320 1432
VGG19 3989 1015
DenseNet121 4133 765
3.5 Temporary learning
One of DL’s shortcomings is the lack of continual learning ability, whereas the brain can learn continuously. The brain can use previous knowledge to incorporate new knowledge, but DL models’ catastrophic forgetting, which deletes previously obtained information to learn the new, prevents continual learning. Since symbols can serve as previous knowledge, we speculate that the temporary mapping between symbols of in-distribution examples and OOD examples may allow a short-term (temporary) learning of OOD examples, which can serve as a stop-gap measure, until a more formal retraining becomes available. To address this possibility, we randomly split NINCO symbols (obtained from 18 classes) into 2 distinct distributions, which are training and test sets. The size of the training and test sets are half of the full size of NINCO symbols. Then, the training symbols were used to construct $CM(i,j)$, and we asked if the test symbols can predict the inputs’ classes. As selecting training/test symbols is done randomly, we resampled them 100 times and reported the results from all 100 experiments (Fig. 7). We observed notable differences between the models but do not know the reasons behind them. More importantly, the accuracy is around 40-50% (Fig. 7) for all 4 models, supporting that internal symbols can be used for short-term learning.
Figure 7: The accuracy of predictions on the OOD examples for ResNet18(Res18), ResNet50(Res50), VGG19 and DenseNet121(DENSE). We resampled trainig and test examples from 18 classes of NINCO dataset [33] for 100 times, and collected their accuracies.
4 Discussion
We probed the existence of internal symbols closely linked to DNNs’ decision making. More specifically, we combined STCert and traditional machine learning algorithms to detect internal symbols. Our analysis of these internal symbols suggest that they can be used to 1) predict levels of accuracies of DNNs’ answers, 2) detect OOD and adversarial inputs and 3) enable temporary learning of OOD examples. Based on our analysis, we propose that decoding of internal symbols could play a crucial role in making DNNs’ operations more reliable and safer.
4.1 Limitations
We should underline that the purpose of our study is not to provide a comprehensive survey of methods that can extract internal symbols and potential utilities of internal symbols. Instead, we focused on searching for positive evidence of the existence of internal symbols associated with DNNs’ decisions and exploring their applications. Below, we list the limitations of our analysis as potential guides for future studies on internal symbols.
  • We selected Mixed_13 and ImageNet models for two reasons. First, if we extract symbols from poorly trained models, the analysis can be biased. Second, as the training set of ImageNet is huge, analyzing them requires substantial computing resources. As Mixed_13 contains images from well-conceived 13 visual categories and balanced number of classes across all categories, it is a good subset to analyze the internal symbols related to ImageNet. However, this raises the question of scalability of our analysis, and we will address this in the future.
  • We extracted symbols on layer-by-layer basis and assumed that they are encoded collectively by neurons with the same receptive fields across channels. However, the feature visualization shows that some single neurons can encode complex features recognizable to humans. We speculate that some symbols are encoded by a set of neurons, and others, by single neurons, but this needs further study.
  • Our analysis searched for symbols related to labels of ImageNet images, but visual objects rarely exist alone in visual scenes. Other objects or background can serve as context to an object of interest. That is, the symbols can interact with one another. Such functional links between symbols are ignored in our analysis.
  • We examined convolutional networks (CNNs) pretrained on ImageNets. Although CNNs are one of the popular architectures, transformers have been widely used, especially in language models. As transformers are based on fully connected layers and self-attention modules, it remains unclear how to probe internal symbols relevant to their decision-making.
4.2 Broader Impacts
We propose a potential algorithm to analyze DL models’ decision. As this algorithm aims to reduce DL models’ mistakes, we expect it could be helpful in making DL more reliable and are not aware of any potential negative impacts.
Appendix A Appendix and Supplementary Material
A.1 STCert implementation
The two versions of STCert were discussed in the original study [20]. The naive STCert uses the bounding boxes returned by the segmentation models as ROIs. The context-aware STCert adds some backgrounds to the bounding boxes to create ROIs. As the context-aware STCert is more robust, we adopted the context-aware STCert in this study. The workflow of STCert can be summarized as follows:
  1. 1.
    Detect bounding boxes.
  2. 2.
    Create 5 candidates of bounding boxes by enlarging ROIs. The first candidate is ROI itself.
  3. 3.
    Crop pixels within the candidates and forward them into DNNs. That is, we have the list of 5 second thought predictions.
  4. 4.
    Examine if the original prediction is included in the list of second predictions.
  5. 5.
    Return the bounding box as ROI, if the original prediction is in the list. If not, ignore the bounding box and return Null.
A.2 Potential links to the brain structure
Our analysis raises the possibility that internal symbols can be used to augment DNNs’ decision-making process. It may even be possible for a second network to use symbols to correct responses in each layer locally during inference. When immediate decisions are necessary, DNNs’ decisions can be used, as they are, but for more reliable answers, a second network may use symbols to modify hidden layers’ responses. This may remind some readers of thalamo-cortical loops of the brain. Our analysis has been inspired by thalamo-cortical loops and indicates that thalamo-cortical loop-like error correction could be constructed using internal symbols to build more reliable DNNs.
Supplementary Table 1: List of classes in the Mixed_13.
Category Class
Fish
‘tench’, ‘goldfish’, ‘great_white_shark’,
‘tiger_shark’, ‘hammerhead’, ‘electric_ray’
Bird
‘cock’, ‘hen’, ‘ostrich’,
‘brambling’, ‘goldfinch’, ‘house_finch’
Dog
‘Chihuahua’, ‘Japanese_spaniel’, ‘Maltese_dog’,
‘Pekinese’, ‘Shih-Tzu’, ‘Blenheim_spaniel’
Carnivore
‘cougar’, ‘lynx’, ‘leopard’,
‘snow_leopard’, ‘jaguar’, ‘lion’
Insect
‘tiger_beetle’, ‘ladybug’,‘ground_beetle’,
‘long-horned_beetle’, ‘leaf_beetle’, ’dung_beetle’
Primate
‘guenon’, ‘patas’, ‘baboon’,
‘macaque’, ‘langur’, ‘colobus’
Car
‘ambulance’, ‘cab’, ‘convertible’,
‘jeep’, ‘limousine’, ‘station_wagon’
Furniture
‘barber chair’, ‘bassinet’, ‘bookcase’,
‘chiffonier’, ‘china_cabinet’, ‘cradle’
Computer
‘desktop_computer’, ‘hand-held_computer’,
‘laptop’, ‘notebook’, ‘slide_rule’, ‘web_site’
Fruit
‘Granny_Smith’, ‘rapeseed’, ‘corn’,
‘acorn’, ‘hip’, ‘buckeye’
Fungus
‘coral_fungus’, ‘agaric’,‘gyromitra’,
‘stinkhorn’, ‘earthstar’, ‘hen-of-the-woods’
Truck
‘fire_engine’, ‘garbage_truck’, ‘moving_van’,
‘pickup’, ‘police_van’, ‘tow_truck’
Supplementary Figure 1: The correlations between the symbols and classes. Top, middle and bottom rows show the correlations observed from ResNet50, VGG19 and DenseNet121. $y$-axis denotes the indices of symbols, $x$-axis denotes the class. (A)-(D), the correlations observed in layers 1-4, respectively.
Supplementary Figure 2: The correlations between symbols and classes in ResNet18. These plots show the correlation maps between 200 consecutive symbols and the label of inputs observed in layer 1 (A), layer 2 (B), layer 3 (C) and layer 4 (D). We note that some consecutive symbols are correlated with the same classes and thus can be merged together, especially in layer 4.